Imagine that the banking system received additional deposits of $100 million and that all the individual banks wish to retain their current liquidity ratio of 20%. If one-half of any additional liquidity is held outside the banking sector, by how much less will deposits have risen compared with (d) above? 


1
Expert's answer
2021-12-12T16:36:55-0500

Solution:

 

Required Reserves = Required Reserve Ratio × Deposits Required Reserves

= 0.20 × 100 = 20 million

Required Reserves = 20 million

 

Additional liquidity held outside = 0.5 x 100 = $50 million

By how much will deposit have risen = 50 – 20 = 30 million